solve the proportion 8/N = 14/24
Answer by algebrahouse.com(1659)

To solve a proportion, cross-multiply:

8/N = 14/24
14N = 8(24) {cross-multiplied}
14N = 192 {multiplied through}
N = 192/14 {divided both sides by 14}
N = 96/7 {reduced}.....you can also divide 96 by 7

For more help from me, visit: www.algebrahouse.com